OpenStudy (anonymous):

Find the break-even point by solving each system using substitution. Show all work. 20. Cost: y = x + 100 Revenue: y = 3x Break-even Point:

OpenStudy (anonymous):

@abb0t any chance you know how to do this?

OpenStudy (abb0t):

This is a system. You are already given \(y\), function. Simply substitute for \(y\) and solve for \(x\)

OpenStudy (abb0t):

\(3x=x+100\)

OpenStudy (anonymous):

oh so 2x=100 x=50

OpenStudy (mathstudent55):

The break even point is where cost = revenue. Since you have y = x + 100 and y = 3x, just set x + 100 = 3x and solve for x.

OpenStudy (abb0t):

Yes

OpenStudy (anonymous):

x=50 but is 50 the break even point?

OpenStudy (anonymous):

@abb0t @mathstudent55

OpenStudy (anonymous):

@mathstudent55